# Representing Sugihara monoids

### v2

@@ 4.1 Closure of $\mathsf{RDInFL}$ under ultraproducts
$\left([\alpha^{-1}_{X}(y)],[x]\right)\in E_{Y}\textnormal{ and }\left([\alpha^{-1}_{X}(y)],[x]\right)\notin\varphi\left([a]\right)$
iff $\left([\alpha^{-1}_{X}(y)],[x]\right)\in\left(\varphi\left([a]\right)\right)^{c}$
iff $\left([x],[\alpha^{-1}_{X}(y)]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
@@
Also observe that $\varphi_{i}(a(i))\cdot^{\mathbf{B}_{i}}\varphi_{i}(c(i))=\left(\varphi(a)\cdot^{\mathbf{B}}\varphi(c)\right)(i)$. Using this in the last step below, we can now see that
$\overline{\varphi}\left([a]\right)\cdot^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\overline{\varphi}\left([c]\right)=\overline{\varphi}\left([a]\right)\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}\overline{\varphi}\left([c]\right)=[\varphi(a)]\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}[\varphi(c)]=\left[\varphi(a)\cdot^{\mathbf{B}}\varphi(c)\right]=\left[\varphi(a\cdot^{\mathbf{A}}c)\right]$
and hence $\overline{\varphi}\left([a]\cdot^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a]\right)\cdot^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\overline{\varphi}\left([c]\right)$. Using the same approach, it is straightforward to show that $\overline{\varphi}$ preserves the operations $\wedge^{\mathbf{A}/\theta_{\mathcal{F}}}$, $\vee^{\mathbf{A}/\theta_{\mathcal{F}}}$, $1^{\mathbf{A}/\theta_{\mathcal{F}}}$ and ${\sim}^{\mathbf{A}/\theta_{\mathcal{F}}}$.
@@
$\overline{\varphi}\left([a]\to^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a\to^{\mathbf{A}}c]\right)$ $=\left[\varphi(a\to^{\mathbf{A}}c)\right]$
$=\left[{\sim}^{\mathbf{B}}\left(-^{\mathbf{B}}\varphi(c)\cdot^{\mathbf{B}}\varphi(a)\right)\right]$
@@ Acknowledgements
The authors are extremely grateful to the two anonymous referees for their very insightful and knowledgeable feedback on the first version on this paper. Their comments enabled us to see that additional results could be possible, and this led to the development of Section 4. The first author would like to thank Chapman University for its hospitality during a research visit in September/October 2023. We would also like to thank Jamie Wannenburg for useful discussions on this topic.

### v3

@@ 4.1 Closure of $\mathsf{RDInFL}$ under ultraproducts
$\left([\alpha^{-1}_{X}(y)],[x]\right)\in E_{Y}\textnormal{ and }\left([\alpha^{-1}_{X}(y)],[x]\right)\notin\varphi\left([a]\right)$ iff $\left([x],[\alpha^{-1}_{X}(y)]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
@@
The proof that $\varphi$ preserves -- is similar, and uses the definition of $\alpha_{Y}$.
@@
Also observe that $\varphi_{i}(a(i))\cdot^{\mathbf{B}_{i}}\varphi_{i}(c(i))=\left(\varphi(a)\cdot^{\mathbf{B}}\varphi(c)\right)(i)$. Using this in the last step below, we can now see that $\overline{\varphi}\left([a]\right)\cdot^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\overline{\varphi}\left([c]\right)=\overline{\varphi}\left([a]\right)\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}\overline{\varphi}\left([c]\right)=[\varphi(a)]\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}[\varphi(c)]=\left[\varphi(a)\cdot^{\mathbf{B}}\varphi(c)\right]=\left[\varphi(a\cdot^{\mathbf{A}}c)\right]$ and hence $\overline{\varphi}\left([a]\cdot^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a]\right)\cdot^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\overline{\varphi}\left([c]\right)$. Using the same approach, it is straightforward to show that $\overline{\varphi}$ preserves the operations $\wedge^{\mathbf{A}/\theta_{\mathcal{F}}}$, $\vee^{\mathbf{A}/\theta_{\mathcal{F}}}$, $1^{\mathbf{A}/\theta_{\mathcal{F}}}$ and ${\sim}^{\mathbf{A}/\theta_{\mathcal{F}}}$.
@@
$\overline{\varphi}\left([a]\to^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a\to^{\mathbf{A}}c]\right)=\left[\varphi(a\to^{\mathbf{A}}c)\right]$ $=\left[{\sim}^{\mathbf{B}}\left(-^{\mathbf{B}}\varphi(c)\cdot^{\mathbf{B}}\varphi(a)\right)\right]$
@@ Acknowledgements
The authors are extremely grateful to the three anonymous referees for their very insightful and knowledgeable feedback on the earlier versions of this paper. Their comments enabled us to see that additional results could be possible, and this led to the development of Section 4. The first author would like to thank Chapman University for its hospitality during a research visit in September/October 2023. We would also like to thank Jamie Wannenburg for useful discussions on this topic.